A ball is dropped from a height of 64 feet. Its height above the earth in feet is given by h(t)=-16t^2+64, where t is the number of seconds after it is dropped.
Find h(1).
How long does it take the ball to fall to the earth?

Answer by stanbon(75887)

A ball is dropped from a height of 64 feet.
Its height above the earth in feet is given by h(t)=-16t^2+64, where t is the number of seconds after it is dropped.
Find h(1) = -16 + 64 = 48 ft
-------------------------------------------
How long does it take the ball to fall to the earth?
The height will be zero at that time.
Solve -16t^2+64 = 0
-16t^2 = -64
t^2 = 4
Positive solution:
t = 2 seconds
=========================
Cheers,
Stan H.